There were 150 tickets sold for a school play. tickets for students were $2.00 and adults $3.00. the total amount of money made was $150. how many more student tickets were sold than adult? what equation(s) couldja make for this? using x and/or y??
Answer by checkley71(8403)

150 tickets for 340 requires x tickets for 2.00 & (150-x) tickets for 3.00 thus
2.00x+3.00(150-x)=340
2x+450-3x=340
-x=340-450
-x=-110
x=110 tickets for $2 then
150-110=40 tickets for $3
proof
110*2+40*3=340
220+120=340
340=340
IN THE FUTURE PRECHECK YOUR PROBLEMS.